# MambaPEFT: Exploring Parameter-Efficient Fine-Tuning for Mamba

### v2

@@ Abstract
An ecosystem of Transformer-based models has been established by building large models with extensive data. Parameter-efficient fine-tuning (PEFT) is a crucial technology for deploying these models to downstream tasks with minimal cost while achieving effective performance. Recently, Mamba, a State Space Model (SSM)-based model, has attracted attention as a potential alternative to Transformers. While many large-scale Mamba-based models have been proposed, efficiently adapting pre-trained Mamba-based models to downstream tasks remains unexplored. In this paper, we conduct an exploratory analysis of PEFT methods for Mamba. We investigate the effectiveness of existing PEFT methods for Transformers when applied to Mamba. We also modify these methods to better align with the Mamba architecture. Additionally, we propose new Mamba-specific PEFT methods that leverage the distinctive structure of Mamba. Our experiments indicate that PEFT performs more effectively for Mamba than Transformers. Lastly, we demonstrate how to effectively combine multiple PEFT methods and provide a framework that outperforms previous works. To ensure reproducibility, we will release the code after publication.
@@ 1 Introduction
To address this fundamental drawback, Gu & Dao (2023) proposed Mamba, a linear-time sequence model that leverages the strengths of State Space Models (SSMs). While Transformers are constructed from attention modules, Mamba is based on the SSM architecture, allowing it to handle long sequences more efficiently. Additionally, Mamba has been shown to perform better than Transformers with the same number of parameters on major tasks such as natural language processing (NLP) (Gu & Dao, 2023) and computer vision (CV) (Zhu et al., 2024). This fact makes Mamba stand out from other post-Transformer architectures with sub-square time complexity (Peng et al., 2023; Sun et al., 2023). Mamba-based models have been proposed across a wide range of domains (Behrouz & Hashemi, 2024; Liang et al., 2024; Zhang et al., 2024b; Li et al., 2024). We believe that Mamba has the potential to go beyond the Transformer ecosystem.
@@ 4.4 Language tasks
In addition to the image tasks, we evaluate our method on language tasks using the vanilla Mamba Gu & Dao (2023). We experiment with a commonsense reasoning task, following the setup and dataset of Hu et al. (2023). This task consists of eight sub-tasks, which are evaluated on each task after training on a dataset that integrates training data from all tasks. We use Pythia (Biderman et al., 2023), which is pre-trained with the same dataset as Mamba, as a Transformer baseline. For a PEFT method for Mamba, we compare with SLL LoRA (Halloran et al., 2024). To the best of our knowledge, SLL LoRA is the only PEFT method for Mamba.
@@ 4.5 Limitations
This paper focuses on an exploratory investigation. Evaluating the applicability of our findings to larger models, such as vision models trained with ImageNet-21K (Ridnik et al., 2021) or large language models with billions of parameters (Touvron et al., 2023), is an area for future work. We will release the code after publication to stimulate future research from communities.
@@ A.7 Experiments with VMamba
A.8 Direction of the Future Work
@@ B.1 Language Tasks
As to the learning rate, we use suitable values for each method. It is set to 5e-5 for the full fine-tuning of Mamba. For SLL LoRA on Mamba, we follow the original learning rate per model size (Halloran et al., 2024). We also use a roughly tuned learning rate for our PEFT methods obtained through the ablation studies in Section A.4. The configurations are shown in Table 7.

### v3

@@ Abstract
An ecosystem of Transformer-based models has been established by building large models with extensive data. Parameter-efficient fine-tuning (PEFT) is a crucial technology for deploying these models to downstream tasks with minimal cost while achieving effective performance. Recently, Mamba, a State Space Model (SSM)-based model, has attracted attention as a potential alternative to Transformers. While many large-scale Mamba-based models have been proposed, efficiently adapting pre-trained Mamba-based models to downstream tasks remains unexplored. In this paper, we conduct an exploratory analysis of PEFT methods for Mamba. We investigate the effectiveness of existing PEFT methods for Transformers when applied to Mamba. We also modify these methods to better align with the Mamba architecture. Additionally, we propose new Mamba-specific PEFT methods that leverage the distinctive structure of Mamba. Our experiments indicate that PEFT performs more effectively for Mamba than Transformers. Lastly, we demonstrate how to effectively combine multiple PEFT methods and provide a framework that outperforms previous works. The source code is available at: https://github.com/sony/mambapeft.
@@ 1 Introduction
To address this fundamental drawback, Gu & Dao (2023) proposed Mamba, a linear-time sequence model that leverages the strengths of State Space Models (SSMs). While Transformers are constructed from attention modules, Mamba is based on the SSM architecture, allowing it to handle long sequences more efficiently. Additionally, Mamba has been shown to perform better than Transformers with the same number of parameters on major tasks such as natural language processing (NLP) (Gu & Dao, 2023) and computer vision (CV) (Zhu et al., 2024). This fact makes Mamba stand out from other post-Transformer architectures with sub-square time complexity (Peng et al., 2023; Sun et al., 2023). Mamba-based models have been proposed across a wide range of domains (Behrouz & Hashemi, 2024; Liang et al., 2024; Zhang et al., 2024b; Li et al., 2024). Mamba has the potential to go beyond the Transformer ecosystem.
@@ 4.4 Language tasks
In addition to the image tasks, we evaluate our method on language tasks using the vanilla Mamba Gu & Dao (2023). We experiment with a commonsense reasoning task, following the setup and dataset of Hu et al. (2023). This task consists of eight sub-tasks, which are evaluated on each task after training on a dataset that integrates training data from all tasks. We use Pythia (Biderman et al., 2023) as the Transformer baseline, which is pre-trained with the same dataset as Mamba. For a PEFT method for Mamba, we compare with SLL LoRA (Halloran et al., 2024). To the best of our knowledge, SLL LoRA is the only PEFT method for Mamba.
@@ 4.5 Limitations
This paper focuses on an exploratory investigation. Evaluating the applicability of our findings to larger models, such as vision models trained with ImageNet-21K (Ridnik et al., 2021) or large language models with billions of parameters (Touvron et al., 2023), is an area for future work. We release the source code to stimulate future research from communities.
@@ 5 Conclusion
Acknowledgements
We would like to thank Takeshi Ohashi, Wei-Yao Wang, Helen Suzuki, and Julian Tanke for their helpful comments to this manuscript.
@@ A.4 Ablation studies on language tasks
Table 6 shows the experimental results of applying LoRA to individual modules in both Mamba and Pythia. In Mamba, applying LoRA to individual modules yielded better performance than SLL LoRA, which applies LoRA to all modules. Conversely, in Pythia, applying LoRA to all modules tended to perform better than applying LoRA to individual modules. These experimental results demonstrate the differences in behavior between Mamba and Transformer when LoRA is applied.
Table 6: Experimental results applying LoRA to individual modules in Pythia and Mamba.
Model Method BoolQ PIQA SIQA HellaSwag WinoGrande ARC-e ARC-c OBQA Avg.
Pythia 160M Full 61.3 62.9 37.1 30.7 50.6 41.5 24.3 27.8 42.0
LoRA 61.0 62.0 36.3 30.3 52.0 38.2 24.6 28.0 41.6
LoRA(query_key_value) 50.4 62.0 36.8 30.2 50.7 40.5 23.9 27.0 40.2
LoRA(dense_4h_to_h) 50.7 61.7 37.9 29.9 49.6 39.7 22.4 27.0 39.8
LoRA(dense_h_to_4h) 55.5 61.3 37.9 29.9 48.7 36.2 22.8 26.8 39.9
LoRA(dense) 49.9 61.6 36.9 30.1 51.1 41.1 23.3 26.0 40.0
[…]
SLL LoRA 56.3 63.3 38.2 34.6 51.6 43.5 23.6 30.6 42.7
Affix-tuning (w/o proj) 55.1 61.4 36.5 32.9 51.5 36.8 23.5 27.2 40.6
LoRA(in_proj) 53.5 62.9 38.2 33.8 53.1 46.4 23.7 30.8 42.8
LoRAp(X) 61.7 64.0 39.5 34.3 52.2 43.5 25.3 29.4 43.7
Pythia 410M Full 55.0 68.4 42.1 40.8 53.9 50.8 26.7 30.0 46.0
LoRA 61.3 67.7 40.8 39.2 54.9 48.1 24.7 28.6 45.7
LoRA(query_key_value) 60.5 67.5 41.3 39.2 54.3 45.4 24.0 29.0 45.1
LoRA(dense_4h_to_h) 61.2 67.1 40.5 39.0 54.5 47.1 25.1 28.6 45.4
LoRA(dense_h_to_4h) 61.5 67.0 39.9 39.0 53.7 46.2 23.8 28.6 45.0
LoRA(dense) 55.7 67.4 40.5 39.8 54.3 45.2 24.1 30.6 44.7
[…]
SLL LoRA 59.5 69.6 42.2 44.1 54.9 50.6 26.3 30.8 47.3
Affix-tuning (w/o proj) 62.0 67.7 39.3 46.3 54.1 47.8 28.2 31.0 47.0
LoRA(in_proj) 55.4 68.6 41.0 44.7 54.1 52.4 28.3 33.4 47.2
LoRAp(X) 60.8 68.8 42.1 44.7 56.2 50.4 27.4 32.2 47.8
Pythia 1B Full 55.0 70.2 42.5 47.5 54.4 54.1 29.7 33.2 48.3
LoRA 60.0 69.3 40.9 45.3 53.6 49.8 27.2 31.0 47.1
LoRA(query_key_value) 58.0 69.4 41.7 44.7 54.1 50.3 27.3 30.8 47.0
LoRA(dense_4h_to_h) 49.2 69.0 42.0 45.2 53.8 52.3 27.5 31.0 46.3
LoRA(dense_h_to_4h) 53.1 68.9 38.9 45.0 52.2 49.7 26.5 30.2 45.6
LoRA(dense) 53.3 69.4 41.0 46.5 52.4 49.4 27.8 32.6 46.6
[…]
SLL LoRA 60.7 72.0 42.4 54.7 56.9 55.3 29.4 34.2 50.7
Affix-tuning (w/o proj) 57.2 71.7 41.4 55.0 55.8 52.6 29.8 33.0 49.6
LoRA(in_proj) 61.7 71.9 44.0 50.8 56.7 56.3 30.5 33.8 50.7
LoRAp(X) 59.9 72.2 44.2 52.8 58.0 53.7 30.8 34.8 50.8
Pythia 1.4B Full 58.6 71.1 42.7 53.6 55.1 58.5 29.9 34.8 50.5
LoRA 60.1 71.3 42.5 50.1 58.9 57.6 29.6 33.6 50.5
LoRA(query_key_value) 60.7 70.9 42.2 50.0 58.3 55.4 28.2 33.2 49.9
LoRA(dense_4h_to_h) 62.2 71.9 41.2 50.5 57.6 55.6 28.8 33.6 50.2
LoRA(dense_h_to_4h) 61.4 70.7 42.9 50.3 57.6 51.7 26.7 32.8 49.3
LoRA(dense) 62.1 71.1 41.6 51.6 58.2 54.7 29.4 31.6 50.0
[…]
SLL LoRA 59.7 73.5 43.1 56.9 60.7 59.7 31.7 36.0 52.7
Affix-tuning (w/o proj) 63.6 74.0 41.9 58.9 60.6 61.6 33.4 36.8 53.9
LoRA(in_proj) 62.6 73.6 43.7 55.6 59.7 58.3 31.7 35.6 52.6
LoRAp(X) 63.1 73.5 42.7 57.7 61.6 60.4 32.9 37.4 53.7
@@ A.7 Experiments with VMamba
A.8 Comparison with Concurrent Work
We discuss the comparison with related studies conducted during the same period in addition to SLL LoRA (Halloran et al., 2024). Galim et al. (2024) proposed SDLoRA, which selectively updates the channels and states of SSM. Unlike SDLoRA, our work emphasizes comprehensive benchmarking and exploration. Furthermore, Galim et al. (2024) pointed out the limitations of expressiveness when applying Prefix-tuning to SSM. In contrast, our proposed Affix-tuning avoids this issue by allowing tokens to be inserted at arbitrary positions. In fact, Affix-tuning shows competitive performance with LoRA on language tasks (table 3).
A.9 Direction of the Future Work
@@ B.1 Language Tasks
As to the learning rate, we use suitable values for each method. For SLL LoRA on Mamba, we follow the original learning rate per model size (Halloran et al., 2024). We also use a roughly tuned learning rate for our PEFT methods obtained through the ablation studies in Section A.4. The configurations are shown in Table 8.